# Bialgebras, and Lie monoid actions in Morse and Floer theory, I

### v1

@@ 3.4. Bimodules of $f$-bialgebras
    ($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}}$) $\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq\begin{cases}\mu^{\widehat{\underline{k}}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\otimes{\alpha}^{1}_{\underline{l}}\text{ if $i$ hits ${\underline{k}}^{l}$,}\\ \mu^{\underline{k}^{l}|\epsilon|\widehat{\underline{k}}^{r}}_{\underline{l}}\otimes{\beta}^{1}_{\underline{l}}\text{ if $i$ hits ${\underline{k}}^{r}$,}\\ \mu^{\widehat{\underline{k}}^{l}|0|\widehat{\underline{k}}^{r}}_{\underline{l}}\otimes\mu^{0|1|0}_{\underline{l}}\text{ if $i$ hits $\epsilon=1$.}\end{cases}$
    In the third case the last (resp. first) entries of $\underline{k}^{l}$ and $\underline{k}^{r}$ are equal to 0, and $\widehat{\underline{k}}^{l}$, $\widehat{\underline{k}}^{r}$ corresponds to removing them.
@@
    (3.65) $(R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b^{l}|\nu|b^{r}}),\ (V^{a^{l}|\epsilon|a^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r},i})$
@@ 3.5. Morphisms of bimodules of $f$-bialgebras
($M^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$ ) $\partial(\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}})=\sum_{GBF_{1}^{f=0}}{\Phi^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}}\circ\mu^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}}+\sum_{GBF_{1}^{f=1}}{\nu^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}\circ\Phi^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}}}.$
And also:
@@
    ($W^{\underline{k}^{l}|1|\underline{k}^{r}}_{b}$ ) $\Phi^{\underline{k}^{l}|1|\underline{k}^{r}}_{\underline{l}}$ $\simeq f^{k^{l}_{1}}_{\underline{l}}\otimes\cdots\otimes f^{k^{l}_{a^{l}}}_{\underline{l}}\otimes\Phi^{k^{l}_{a^{l}+1}|1|k^{r}_{0}}_{\underline{l}}\otimes g^{k^{r}_{1}}_{\underline{l}}\otimes\cdots\otimes g^{k^{r}_{a^{r}}}_{\underline{l}},$
    and
    ($X^{0|1|0}_{b}$ ) $\Phi^{0|1|0}_{\underline{l}}\simeq\Phi^{0|1|0}_{1}\otimes\cdots\otimes\Phi^{0|1|0}_{1}.$
@@
    ($W^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$ ) $\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{1}}\otimes\cdots\otimes\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{b}},$
    and
    ($X^{a^{l}|\epsilon|a^{r}}_{1}$ ) $\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{1}\simeq f^{1}_{1}\otimes\cdots\otimes f^{1}_{1}\otimes\Phi^{0|\epsilon|0}_{1}\otimes g^{1}_{1}\otimes\cdots\otimes g^{1}_{1},$
    with $\Phi^{0|\epsilon|0}_{1}=1$ if $\epsilon=0$.
@@ 3.6. Bimodule categories and functors
For $\epsilon=1$, one can think of $\mu^{\underline{k}^{l}|1|\underline{k}^{r}}_{\underline{L}}$ as a rectangular box as in Figure 18. For $\epsilon=0$ they only depend on $\underline{L}$ through $\underline{l}$, so we also denote them $\mu^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{l}}$. We may also do so when $\epsilon=1$, although this is an abuse of notation (as for $A_{\infty}$-categories). These operations are required to be such that $\mu^{0|1|0}_{(L)}=id_{\mathcal{M}(L,L)}$, and to satisfy the obvious analogs
$(R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{L}}),\ (E^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{L}}),\ (V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b}),\ (V^{a^{l}|\epsilon|a^{r}}_{\underline{L}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{L}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{L},i})$
of equations ($R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$), ($E^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{l}}$), ($V^{\underline{k}^{l}|1|\underline{k}^{r}}_{b}$), ($V^{\underline{k}^{l}|0|\underline{k}^{r}}_{b}$),( $V^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$), ($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}}$), ($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l},i}$), i.e. replacing multi-indices $\underline{l},\underline{l}^{0},\underline{l}^{1}$ by decorated multi-indices $\underline{L},\underline{L}^{0},\underline{L}^{1}$.
@@
satisfying the obvious analogs of the (M), (W), (X) relations of Definition 3.11.
@@ 3.7. $d$-categories, bimodules and functors
$(R(\mathbb{L})^{\underline{k}}_{\underline{l}}),\ (V(\mathbb{L})^{\underline{k}}_{b}),\ (V(\mathbb{L})^{a}_{\underline{l}}),\ (D(\mathbb{L})^{\underline{k};i}_{\underline{l}}),\ (D(\mathbb{L})^{\underline{k}}_{\underline{l};i}),$
as well as $\alpha(L)^{1}_{1}=id_{\mathcal{A}(L,L)}$.

### v2

@@ 3.4. Bimodules of $f$-bialgebras
    ($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}}$) $\mu^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq\begin{cases}\mu^{\widehat{\underline{k}}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\otimes\widetilde{\alpha}^{1}_{\underline{l}}\text{ if $i$ hits ${\underline{k}}^{l}$,}\\ \mu^{\underline{k}^{l}|\epsilon|\widehat{\underline{k}}^{r}}_{\underline{l}}\otimes\widetilde{\beta}^{1}_{\underline{l}}\text{ if $i$ hits ${\underline{k}}^{r}$,}\\ \mu^{\widehat{\underline{k}}^{l}|0|\widehat{\underline{k}}^{r}}_{\underline{l}}\otimes\widetilde{\mu}^{0|1|0}_{\underline{l}}\text{ if $i$ hits $\epsilon=1$,}\end{cases}$
    where $\widetilde{\alpha}^{1}_{\underline{l}}$ and $\widetilde{\beta}^{1}_{\underline{l}}$ are as in Definition 3.2, and $\widetilde{\mu}^{0|1|0}_{\underline{l}}=\widetilde{\mu}^{0|1|0}_{l_{1}}\otimes\cdots\otimes\widetilde{\mu}^{0|1|0}_{l_{b}}$ is defined analogously, with $\widetilde{\mu}^{0|1|0}_{1}=id_{M}$ and $\widetilde{\mu}^{0|1|0}_{2}={\mu}^{0|1|0}_{2}$. In the third case the last (resp. first) entries of $\underline{k}^{l}$ and $\underline{k}^{r}$ are equal to 0, and $\widehat{\underline{k}}^{l}$, $\widehat{\underline{k}}^{r}$ corresponds to removing them.
@@
    (3.54) $(T^{a^{l}|\epsilon|a^{r}}_{b^{l}|\nu|b^{r}}),\ (R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b^{l}|\nu|b^{r}}),\ (V^{a^{l}|\epsilon|a^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}^{l}|\nu|\underline{l}^{r}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}^{l}|\nu|\underline{l}^{r},i})$
@@ 3.5. Morphisms of bimodules of $f$-bialgebras
($M^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$ ) $0=\sum{(-1)^{\rho_{0}}\cdot\Phi^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}\circ\mu^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}}+(-1)^{\rho_{1}}\cdot\nu^{\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0}}_{\underline{l}_{0}}\circ\Phi^{\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1}}_{\underline{l}_{1}}},$
where the sum runs over $(\underline{k}^{l}|\epsilon|\underline{k}^{r})=(\underline{k}^{l}_{1}|\epsilon|\underline{k}^{r}_{1})\sharp(\underline{k}^{l}_{0}|\epsilon|\underline{k}^{r}_{0})$ and $\underline{l}=\underline{l}^{0}\sharp\underline{l}^{1}$. And also:
@@
    ($W^{\underline{k}^{l}|1|\underline{k}^{r}}_{b}$ ) $\Phi^{\underline{k}^{l}|1|\underline{k}^{r}}_{\underline{l}}$ $\simeq f^{k^{l}_{1}}_{\underline{l}}\otimes\cdots\otimes f^{k^{l}_{a^{l}}}_{\underline{l}}\otimes\Phi^{k^{l}_{a^{l}+1}|1|k^{r}_{0}}_{\underline{l}}\otimes g^{k^{r}_{1}}_{\underline{l}}\otimes\cdots\otimes g^{k^{r}_{a^{r}}}_{\underline{l}}.$
@@
    ($W^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$ ) $\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}\simeq\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{1}}\otimes\cdots\otimes\Phi^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{l_{b}}.$
@@ 3.6. Bimodule categories and functors
For $\epsilon=1$, one can think of $\mu^{\underline{k}^{l}|1|\underline{k}^{r}}_{\underline{L}}$ as a rectangular box as in Figure 15. For $\epsilon=0$ they only depend on $\underline{L}$ through $\underline{l}$, so we also denote them $\mu^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{l}}$. We may also do so when $\epsilon=1$, although this is an abuse of notation (as for $A_{\infty}$-categories). These operations are required to satisfy the obvious analogs
$(R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{L}}),\ (T^{a^{l}|\epsilon|a^{r}}_{\underline{L}}),\ (V^{a^{l}|\epsilon|a^{r}}_{\underline{L}}),\ (V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b}),\ (E^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{L}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{L}}),\ (D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{L},i})$
of equations ($R^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l}}$), $\left(T^{a^{l}|\epsilon|a^{r}}_{b}\right)$, ( $V^{a^{l}|\epsilon|a^{r}}_{\underline{l}}$), ($V^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{b}$), ($E^{\underline{k}^{l}|0|\underline{k}^{r}}_{\underline{l}}$),($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r},i}_{\underline{l}}$), ($D^{\underline{k}^{l}|\epsilon|\underline{k}^{r}}_{\underline{l},i}$), i.e. replacing multi-indices $\underline{l},\underline{l}^{0},\underline{l}^{1}$ by decorated multi-indices $\underline{L},\underline{L}^{0},\underline{L}^{1}$.
@@
satisfying the obvious analogs of the (M) and (W) relations of Definition 3.15.
@@ 3.7. $d$-categories, bimodules and functors
$(R(\mathbb{L})^{\underline{k}}_{\underline{l}}),\ (T(\mathbb{L})^{a}_{b}),\ (V(\mathbb{L})^{\underline{k}}_{b}),\ (V(\mathbb{L})^{a}_{\underline{l}}),\ (D(\mathbb{L})^{\underline{k};i}_{\underline{l}}),\ (D(\mathbb{L})^{\underline{k}}_{\underline{l};i}).$